OpenStudy (sboodhoo):

What is the simplified expression for 3^3x3^3/3^4? 3^0 3^1 3^2 3^4

OpenStudy (anonymous):

hmmm maybe justtt maybe "B" but i need support so @Mexican_White_boy

OpenStudy (anonymous):

cause i'm just saying the first half would equal up to 729 over 81 and simplified would be 9/1 lol wait so it would be 3^2

OpenStudy (anonymous):

cause 9/1 would be 9 and the only thing here that equals 9 is 3^2 lol im so slow

OpenStudy (er.mohd.amir):

3^6/3^4=3^2=9

OpenStudy (sboodhoo):

A closed container has 3.06 ⋅ 10^23 atoms of a gas. Each atom of the gas weighs 1.67 ⋅ 10^−24 grams. Which of the following shows and explains the approximate total mass, in grams, of all the atoms of the gas in the container? 0.47 grams, because (3.06 + 1.67) ⋅ (10^23 ⋅ 10^−24) = 4.73 ⋅ 10^−1 0.51 grams, because (3.06 ⋅ 1.67) ⋅(10^23 ⋅ 10^-24) = 5.1102 ⋅ 10^-1 4.73 grams, because (3.06 + 1.67) ⋅ (10^23 ⋅ 10^−24) = 4.73 5.11 grams, because (3.06 ⋅ 1.67) ⋅ (10^23 ⋅ 10^−24) = 5.1102